what is the area of a 2ft wide walk around the outside of a garden that is 30ft long and 20ft wide?
Answer by Nate(3500)

width = 2 + 20 + 2 = 24
length = 2 + 30 + 2 = 34
area of whole area = 24*34
area of garden = 20*30
area of walk = 24*34 - 20*30
area of walk = 816 - 600 = 216ft^2